Kelly Company has a current ratio greater than 1 and an acid-test ratio less than 1. how would cash payments to suppliers to reduce accounts payable affect these ratios?

current ratio quick ratio
a. decreased decreased
b. decreased increased
c. increased decreased
d. increased increased
Answer #1
Current Ratio=Current Assets (Includes Cash)/Current Liabilites
Acid Test Ratio=Quick Assets (Includes Cash exculdes Stock and Prepaid Liabilties)/Current Liabilites
Taking an Example:
Current Ratio is Greater than 1:
Assumed CA= 150000 , CL=100000
Current Ratio= 150000/100000
Current Ratio= 1.5
Suppose 50000 is Paid to settle Accounts Payable
Assumed CA= (150000-50000) , CL=(100000-50000)
Current Ratio= 100000/50000
Current Ratio= 2.0
Current Ratio will increase
Quick Ratio is Less than 1:
Assumed CA= 80000 , CL=100000
Current Ratio= 80000/100000
Quick Ratio= 0.8
Suppose 50000 is Paid to settle Accounts Payable
Assumed CA= (80000-50000) , CL=(100000-50000)
Current Ratio= 30000/50000
Qucik Ratio= 0.6
Quick Ratio will Decrease
Answer is C